OpenStudy (anonymous):

What two numbers have a sum of -5 and a product of 6?

OpenStudy (anonymous):

let "a" be the value of the first number let "b" be the value of the 2nd number sum means add product means multiply the sum of 2 numbers is -5 a+b=-5 the product of 2 numbers is 6 a*b=6 solve by using substitution

OpenStudy (anonymous):

yeah i got that but i cant figure out how to solve it

OpenStudy (anonymous):

lets say i solve for a for the product of 2 numbers a*b=6 solving for a a=6/b now since i know "a", i can substitute that into the other equation

OpenStudy (anonymous):

yeap so 6/b+b=-5

OpenStudy (anonymous):

What two numbers can you think of at the top of your head that can be used to equal 6 & 5? :)

OpenStudy (anonymous):

I'd rather them learn how to solve it than just be given the answer and not know how to solve it..

OpenStudy (anonymous):

well yea, you could just guess and check, but there's still a process and dont give out answers, you know better than that

OpenStudy (anonymous):

thanks for the answer, @completeidiot was just trying to explain how to solve it and show working out, no need for the "common sense" thanks

OpenStudy (anonymous):

yes a quicker method would just be to determine all the multiples of -6, then just plug in the values to see which works

OpenStudy (anonymous):

okay thankyou very much :)

OpenStudy (anonymous):

im just gonna go through all the steps a=6/b a+b=-5 6/b+b=-5 6+b^2=-5b b^2+5b+6=0 (b+2)(b+3)=0 b= -2, b=-3